Alternative Exercise 3.119 A pistol that fires a signal flare gives the flare an initial speed (muzzle speed) of 104 m/s Part A If the flare is fired at an angle of 61.0°aboe the horizontal on the level salt flats of Utah, what is its horizontal range? You can ignore air resistance. ANSWER km Part B If the flare is fired at the same angle over the flat Sea of Tranquility on the moon, where g-1.6m/s, what is its horizontal range? ANSWER: km Problem 3.61 A boy 11.0 m above the ground in a tree throws a ball for his dog, who is standing right below the tree and starts running the instant the ball is thrown. If the boy throws the ball upward at 45.0 above the horizontal, at 7.50 m/s Part A How fast must the dog run to catch the ball just as it reaches the ground? ANSWER: m/s Part B How far from the tree will the dog catch the ball? ANSWER:
Answer #1

1)

A)

For the vertical component of the motion :

s = ut + 0.5*at^2

So, 0 = 104*sin(61 deg)*t + 0.5*(-9.8)*t^2

So, t = 18.6 s

So, horizontal range,

R = 104*cos(61 deg)*18.6 = 938 m = 0.938 km <----- answer

B)

Here ,

0 = 104*sin(61 deg)*t + 0.5*(-1.6)*t^2

So, t = 113.7 s

So, R = 104*cos(61 deg)*113.7 = 5733 m <-------- answer

2)

A)

For vertical component :

s = ut + 0.5*at^2

So, -11 = 7.5*sin(45 deg)*t + 0.5*(-9.8)*t^2

So, t = 2.13 s

So, horizontal distance traveled = 7.5*cos(45 deg)*2.13 = 11.3 m

So, speed of the dog = 11.3/2.13 = 5.3 m/s <----------- answer

B)

Horizontal Distance , d = 11.3 m <--------- answer (solved above)